OpenStudy (lieutenantgeneral):

What is the circumference of this circle? Use π ≈ 3.14. Round your answer to the nearest tenth of a centimeter. Diamter equals 15cm

OpenStudy (johnweldon1993):

Circumference = 2pi * r R = radius = Half the diameter...so Diameter = 15cm so half that is 7.5cm Now just plug that into the formula above \[\large \text{circumference} = 2\pi*7.5cm\]

OpenStudy (lieutenantgeneral):

Thanks

OpenStudy (lieutenantgeneral):

A ramp leads up to a building. The top of the ramp is 5 feet above the ground, and the bottom is 15 feet from the building as shown. What is the length of the ramp? Round your answer to the nearest tenth of a foot.

OpenStudy (lieutenantgeneral):

height 5cm lenght 15cm

OpenStudy (anonymous):

no it's \[2 \pi r ^{2}\]

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

@LadyKatsaTheGraceling Circumference = 2pi * radius Area = pi * r^2 Somehow you are combining the two equations

OpenStudy (lieutenantgeneral):

don't combine those are 2 different questions

OpenStudy (anonymous):

Well according to the math website I got the formula from…

OpenStudy (johnweldon1993):

@LieutenantGeneral You would use Pythagorean theorem here The Hypotenuse (length here) is equal to \[\large \sqrt{5^2 + 15^2}\]

OpenStudy (johnweldon1993):

Please post that line @LadyKatsaTheGraceling so I may see...

OpenStudy (johnweldon1993):

link* sorry

OpenStudy (lieutenantgeneral):

i about to close the question

OpenStudy (anonymous):

https://xkcd.com/1184/